Determine the km and vmax from this data.

1/[S]
0.1
0.033333333
0.02
0.01
0.006666667
0.002
1/V0
34.3
31.95
27.2
25.51
20.16
19.6
Answer #1

1/V0 = (Km/Vm[S])+1/Vm

y = mx+c

1/[s]

1/V0

0.1

34.3

0.033

31.95

0.02

27.2

0.01

25.51

0.0066

20.16

0.002

19.6

comparing from the chat value of slope and intercept

m = 136.18 = Km/Vm

c = 22.56 = 1/Vm

so Vm = 0.044

and Km = 6